# A Critical Study of Automatic Evaluation in Sign Language Translation

### v1

@@ 1.  Introduction
Sign languages are the primary communication systems for millions of deaf and hard-of-hearing people across the globe. They are highly expressive visual languages that convey meaning through hand signs, facial expressions, mouthings, and body posture (Stokoe, 1980). Automatic sign language translation (SLT) aims to translate sign language videos into spoken language text to help bridge the communication gap between the deaf and hearing communities (Camgoz et al., 2018). However, evaluating SLT systems remains a significant challenge, since current evaluation practices rely solely on textual outputs, offering little insight into how effectively the visual-linguistic nature of sign language is captured. Human evaluation, though often the most reliable and serving as the gold standard in evaluating SLT, is costly and time-consuming, making it difficult to scale. As a result, current SLT models typically rely on text-based automatic evaluation metrics. Consequently, a valid measure of translation quality that accounts for this cross-modal grounding is still missing in SLT (Müller et al., 2023). In addition to the limitations of current evaluation metrics, SLT systems are also prone to hallucinations (Zhang et al., 2023; Hamidullah et al., 2025a), fluent but incorrect translations that misrepresent the signed input. Evaluating how metrics respond to such cases is crucial for understanding their reliability.
When translating from sign language to spoken language text, traditional lexical overlap machine translation (MT) metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) are commonly adopted. However, these text-based metrics fall short in capturing the multimodal complexity of sign language. To address this gap, recent studies have proposed evaluation metrics tailored to sign language tasks. SignBLEU (Kim et al., 2024), for instance, operates at the gloss level and aligns better with human judgments, while SiLVERScore (Imai et al., 2025) introduces an embedding-based approach for assessing sign language generation. Yet, SignBLEU depends on gloss annotations, and SiLVERScore has not been evaluated in the context of SLT. Beyond these, embedding-based and large language model (LLM)-based metrics developed for text-to-text MT—such as BLEURT (Sellam et al., 2020) and GEMBA (Kocmi and Federmann, 2023) have shown strong correlations with human judgments in text-to-text MT (Freitag et al., 2024). Nevertheless, these approaches remain underexplored in SLT, and there is still no comprehensive analysis of their strengths and limitations for evaluating SLT systems.
@@ 2.1.  Evaluation Metrics
SLT models have traditionally relied on lexical overlap MT metrics for automatic translation. Specifically, BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) have been widely used following the neural SLT work by Camgoz et al. (2018). Following Müller et al. (2022), studies in SLT have adopted the embedding-based metric BLEURT (Sellam et al., 2020). However, these metrics do not take the visual source input into account. Recently, Kim et al. (2024) proposed SignBLEU, an evaluation metric designed to reduce information loss when assessing SLT outputs. By operating at the gloss level, it achieves stronger alignment with human judgments than prior text-based metrics. Nevertheless, SignBLEU remains constrained to the text modality and relies on manually curated gloss annotations, which limits scalability. More recently, Imai et al. (2025) introduced SiLVERScore, an embedding-based metric that evaluates sign language generation in a multimodal semantic space. While it shows promising results, SiLVERScore was not evaluated on SLT.
The progress and advancements in LLMs have motivated research into their potential applications for automated assessment across various tasks. Liu et al. (2023) pioneered LLM-based automated assessment by proposing G-Eval, a framework that uses Chain-of-Thought reasoning (Wei et al., 2022) along with task descriptions and evaluation metrics to evaluate performance on tasks including text summarization and dialogue generation. Following this direction, Kocmi and Federmann (2023) proposed GEMBA, a GPT-based evaluation metric to assess the MT quality both with- and without reference in a zero-shot prompting fashion. Recently, inspired by G-Eval, Tong et al. (2025) extended the framework to image and video caption evaluation using GPT-4o, achieving a high correlation with human judgments.
@@
SLT aims to convert sign language videos into spoken language text. SLT models are often categorized into gloss-free and methods with gloss supervision. Camgoz et al. (2018) pioneered SLT by framing it as a neural machine translation (NMT) task, and further extending it by leveraging transformers for end-to-end SLT (Camgoz et al., 2020). MSKA-SLT (Guan et al., 2025) sets a new state-of-the-art on the Phoenix-2014T (Camgoz et al., 2018) and CSL-Daily (Zhou et al., 2021) benchmarks as a gloss-based model, with BLEU scores of $29.03$ and $25.52$, respectively. The approach models keypoints extracted from the face, body, and both hands using an attention mechanism, and is jointly optimized with a CTC loss and a self-distillation loss. However, glosses require manual human annotation, which is labor-intensive and difficult to scale. In contrast, recent approaches attempt to model SLT at the sentence level and a gloss-free fashion. Building on this objective, recent works have incorporated LLMs into SLT frameworks (Gong et al., 2024; Wong et al., 2024; Chen et al., 2024), employed contrastive learning objectives (Cheng et al., 2023), replaced gloss supervision with sentence-level embeddings (Hamidullah et al., 2024), and extended SLT to multilingual scenarios (Yin et al., 2022; Yazdani et al., 2025; Tan et al., 2025; Hamidullah et al., 2025b).
@@ 2.3.  Hallucination
A model is said to hallucinate when it generates content that either lacks logical coherence or diverges from the source material, asserting information or events that are unfounded or inconsistent with the known facts (Ji et al., 2023). While hallucination detection has received increasing attention in the MT community, with studies on multilingual comparisons (Guerreiro et al., 2023) and LLM-based approaches for both low- and high-resource languages (Benkirane et al., 2024), research on hallucination in SLT remains limited. Zhang et al. (2023) reported, based on manual analysis, that SLT models such as SLTUNET suffer greatly from hallucinations, where the generated translations often show limited correlation with the sign video. More recently, Hamidullah et al. (2025a) proposed a hallucination detection framework and further demonstrated that gloss-free SLT models are particularly prone to severe hallucinations compared to their gloss-based counterparts. These findings highlight the need for a deeper understanding of how evaluation metrics respond to hallucinated outputs and how robust current SLT models are under such conditions.
@@ 3.1.  Task and Datasets
In our experiments, we evaluate the limitations of current SLT metrics with respect to paraphrasing (including word reordering), hallucination sensitivity, and variation in translation length. We use the Phoenix-2014T  [^1] Phoenix-2014T was selected because it is the most widely used benchmark in SLT research and contains gloss-level annotations necessary for evaluating gloss-based models. dataset (Camgoz et al., 2018), covering weather forecasts in German sign language (DGS) to evaluate and compare the performance of various models in the above aspects.
@@
SLT models are prone to hallucinations, where the generated text diverges from the visual input. Yet, there has been no systematic investigation into how such hallucinations affect translation quality or the reliability of automatic SLT evaluation metrics. Motivated by findings from Benkirane et al. (2024), which demonstrate the effectiveness of LLMs for hallucination detection in MT, we employ Llama-3-70B to identify hallucinated outputs in a reference-based setting. [^2] The prompt used for hallucination detection is available in Appendix A.1. To analyze how evaluation metrics behave under varying degrees of hallucination, we consider a Severity Ranking scheme that categorizes outputs into four levels: No Hallucination, Small Hallucination, Partial Hallucination, and Full Hallucination.
@@ 3.2.  Evaluation Metrics
For our analysis, we consider the metrics as per Müller et al. (2022, 2023), specifically BLEU [^3] BLEU|nrefs:1|bs:25|tok:none|eff:no| case:mixed|smooth:exp|version:2.5.1 (via SacreBLEU; Post, 2018) for lexical overlap, ROUGE (Lin, 2004) [^4] ROUGE|metrics:rougeL|nrefs:1| stemmer:true|bs:25|version:0.1.2 for recall-oriented n-gram overlap, chrF (Popović, 2015)555chrF|nrefs:1|bs:25|case:mixed|eff:yes| nc:6|nw:2|space:no|version:2.5.1 that uses character $n$-grams, and embedding-based BLEURT (Sellam et al., 2020) [^6] BLEURT v0.0.2 using checkpoint BLEURT-20. for semantic quality. We use bootstrap resampling for statistical significance reporting.
We also extend the traditional range of SLT evaluation metrics by presenting a sign language adapted version of G-Eval (Liu et al., 2023). G-Eval introduces a framework that leverages prompting techniques to generate evaluation scores that strongly correlate with human preferences. The prompt is organized into three modules: (1) evaluation dimensions, which specify the aspects to be judged; (2) step-by-step reasoning, where Chain-of-Thought (CoT) (Wei et al., 2022) guides the LLM through the evaluation procedure; and (3) scoring with references, which constrains the output format and incorporates human translations as ground truth. Following Sato et al. (2024), we prompt the LLMs to evaluate the translation quality of the models in Section 3.3 for Adequacy and Fluency, each rated on a 5-point Likert scale (1155). Following Fomicheva et al. (2022); Kocmi and Federmann (2023), we further include GEMBA zero-shot Direct Assessment (Kocmi and Federmann, 2023), where the LLM is prompted to evaluate each translation hypothesis and provide a quality score ranging from 0 (completely incorrect) to $100$ (perfect translation). Our LLM experimental setup includes four models: GPT-4.1-nano, Qwen3-8B, Llama-3.1-8B, and QWQ-32B. These were chosen to include both a proprietary LLM and also the open-source ones. [^7] All prompts used for G-Eval and GEMBA are provided in Appendices A.2 and A.3.
@@
SLT models are either gloss-based or gloss-free. We include three gloss-free models: SpaMo (Hwang et al., 2025), SEM-SLT (Hamidullah et al., 2024), and Signformer (Yang, 2024). We also include TwoStream-SLT (Chen et al., 2022b) as a gloss-based model. We retrain these models to reproduce their reported results when the original model checkpoints are unavailable; however, our reproduced scores differ slightly from those originally reported.
@@
With the goal of edge AI, Signformer uses a combination of novel convolution, attention, and positional encoding to achieve competitive performance compared to gloss-based models while being quite smaller. Since the original paper did not specify how video features were extracted, we instead employ the S3D model pre-trained on both WLASL for ASL word recognition (Li et al., 2020) and the Kinetics-400 dataset for human action recognition (Kay et al., 2017), following the approach by (Chen et al., 2022a). We set learning rate to 0.001 and train it for a maximum of 100 epochs.
@@ 4.1.  Impact of Paraphrasing on SLT Evaluation
We present the results of LLM-based evaluators, G-Eval (Adequacy and Fluency) and GEMBA, comparing paraphrased and original translations on the Phoenix-2014T test set in Table 2. Most notably, paraphrasing generally improves G-Eval Adequacy and Fluency and GEMBA scores, consistent with the pattern observed for BLEURT. Llama3.1:8b is the only model that rates paraphrased translations as less fluent than the original ones. Among all evaluated LLMs, Llama3.1:8b tends to assign the highest Adequacy and Fluency scores, whereas gpt-4.1-nano assigns the lowest. We observe that LLM-based evaluators tend to prefer paraphrased translations over the original ones. This tendency likely arises because the paraphrases are themselves generated by an LLM, which may introduce stylistic patterns or linguistic preferences that align with the evaluator’s own training distribution, thereby biasing its judgments. Interestingly, this observation aligns with previous studies suggesting that LLMs exhibit a bias toward LLM-generated or paraphrased text (Xu et al., 2024; Liu et al., 2023). To be able to make more general statements about evaluation metrics, we show the Pearson correlation between the lexical metrics, BLEURT, and GEMBA, computed using the average scores from Table 2 in Figure 1. We notice a high correlation among lexical overlap metrics, and GEMBA is also highly correlated with BLEURT, as both incorporate learned language model representations in their scoring. Our previous observations from Tables 1 and 2 are confirmed, with lexical metrics (BLEU, chrF, ROUGE) yielding consistent rankings, while LLM-based evaluations align closely with BLEURT, ranking TwoStream-SLT highest, followed by SpaMo, SEM-SLT, and Signformer.
@@ 4.2.  Impact of Hallucination on SLT Evaluation
We first analyze how hallucinations impact automatic metrics and then examine how robust SLT models are in the case of hallucination. Figure 2 illustrates the impact of hallucinations, categorized by their severity ranking, on various evaluation metrics across four SLT models. We omitted G-Eval in this setting because its results are reported on a Likert scale.  [^8] Typically ranging from 1 (strongly disagree/poor) to 5 (strongly agree/excellent) (Likert, 1932). Moreover, we report hallucination detection results for GEMBA using gpt-4.1-nano. Notably, results show that there is a monotonic decline across all metrics, meaning that as hallucination severity level increases, translation quality drops. However, BLEU is an extremely sensitive metric; in cases such as Partial Hallucinations or Full Hallucinations, even minor rephrasings that remain semantically correct can cause a sharp decrease in BLEU scores. ROUGE and chrF decline more smoothly, since they account for partial overlaps at the subword or character level, respectively. GEMBA and BLEURT follow a similar trend, as they can still capture some semantic meaning even in extreme Full Hallucination cases where the surface wording differs substantially. However, we suspect that these metrics may under-penalize subtle hallucinations when the generated text is fluent, thereby overestimating the overall score. Similar observations regarding BLEURT’s overestimation and the data leakage issues in LLMs have been discussed by Zeng et al. (2024).
@@
Different from the main focus of the paper that concentrates on evaluating evaluation metrics, here we briefly take the opportunity to investigate the behavior of SLT models when hallucinations occurs. Overall, TwoStream-SLT and SpaMo demonstrate a consistent trend across all evaluation metrics, showing less hallucinations compared to other models. In contrast, Signformer exhibits notably lower average performance, aligning with the results presented in Table 1. To analyze the extent to which SLT models produce hallucinations, we present in Table 3 a breakdown of the number of hallucinated translations according to severity level. The results reveal that gloss-based TwoStream-SLT yields the lowest Full Hallucinations (24), indicating higher robustness to severe errors, whereas Signformer produces the highest number (177), confirming our earlier observations regarding its overall vulnerability to hallucination. This finding is in line with previous work showing that gloss-free models exhibit higher hallucination rates than gloss-based models (Hamidullah et al., 2025a).

### v2

@@ 1.  Introduction
Sign languages are the primary communication systems for millions of deaf and hard-of-hearing people across the globe. They are highly expressive visual languages that convey meaning through hand signs, facial expressions, mouthings, and body posture (stokoe1980sign). Automatic sign language translation (SLT) aims to translate sign language videos into spoken language text to help bridge the communication gap between the deaf and hearing communities (Camgoz_2018_CVPR). However, evaluating SLT systems remains a significant challenge, since current evaluation practices rely solely on textual outputs, offering little insight into how effectively the visual-linguistic nature of sign language is captured. Human evaluation, though often the most reliable and serving as the gold standard in evaluating SLT, is costly and time-consuming, making it difficult to scale. As a result, current SLT models typically rely on text-based automatic evaluation metrics. Consequently, a valid measure of translation quality that accounts for this cross-modal grounding is still missing in SLT (muller-etal-2023-findings). In addition to the limitations of current evaluation metrics, SLT systems are also prone to hallucinations (zhang2023sltunet; hamidullah2025grounding), fluent but incorrect translations that misrepresent the signed input. Evaluating how metrics respond to such cases is crucial for understanding their reliability.
When translating from sign language to spoken language text, traditional lexical overlap machine translation (MT) metrics such as BLEU (papineni-etal-2002-bleu) and ROUGE (lin-2004-rouge) are commonly adopted. However, these text-based metrics fall short in capturing the multimodal complexity of sign language. To address this gap, recent studies have proposed evaluation metrics tailored to sign language tasks. SignBLEU (kim-etal-2024-signbleu), for instance, operates at the gloss level and aligns better with human judgments, while SiLVERScore (imai-EtAl:2025:RANLP2) introduces an embedding-based approach for assessing sign language generation. Yet, SignBLEU depends on gloss annotations, and SiLVERScore has not been evaluated in the context of SLT. Beyond these, embedding-based and large language model (LLM)-based metrics developed for text-to-text MT—such as BLEURT (sellam-etal-2020-bleurt) and GEMBA (kocmi-federmann-2023-gemba) have shown strong correlations with human judgments in text-to-text MT (freitag-etal-2024-llms). Nevertheless, these approaches remain underexplored in SLT, and there is still no comprehensive analysis of their strengths and limitations for evaluating SLT systems.
@@ 2.1.  Evaluation Metrics
SLT models have traditionally relied on lexical overlap MT metrics for automatic translation. Specifically, BLEU (papineni-etal-2002-bleu) and ROUGE (lin-2004-rouge) have been widely used following the neural SLT work by Camgoz_2018_CVPR. Following muller-etal-2022-findings, studies in SLT have adopted the embedding-based metric BLEURT (sellam-etal-2020-bleurt). However, these metrics do not take the visual source input into account. Recently, kim-etal-2024-signbleu proposed SignBLEU, an evaluation metric designed to reduce information loss when assessing SLT outputs. By operating at the gloss level, it achieves stronger alignment with human judgments than prior text-based metrics. Nevertheless, SignBLEU remains constrained to the text modality and relies on manually curated gloss annotations, which limits scalability. More recently, imai-EtAl:2025:RANLP2 introduced SiLVERScore, an embedding-based metric that evaluates sign language generation in a multimodal semantic space. While it shows promising results, SiLVERScore was not evaluated on SLT.
The progress and advancements in LLMs have motivated research into their potential applications for automated assessment across various tasks. liu-etal-2023-g pioneered LLM-based automated assessment by proposing G-Eval, a framework that uses Chain-of-Thought reasoning (10.5555/3600270.3602070) along with task descriptions and evaluation metrics to evaluate performance on tasks including text summarization and dialogue generation. Following this direction, kocmi-federmann-2023-gemba proposed GEMBA, a GPT-based evaluation metric to assess the MT quality both with- and without reference in a zero-shot prompting fashion. Recently, inspired by G-Eval, Tong_He_Shao_Yeung_2025 extended the framework to image and video caption evaluation using GPT-4o, achieving a high correlation with human judgments.
@@
SLT aims to convert sign language videos into spoken language text. SLT models are often categorized into gloss-free and methods with gloss supervision. Camgoz_2018_CVPR pioneered SLT by framing it as a neural machine translation (NMT) task, and further extending it by leveraging transformers for end-to-end SLT (camgoz2020sign). MSKA-SLT (10.1016/j.patcog.2025.111602) sets a new state-of-the-art on the Phoenix-2014T (Camgoz_2018_CVPR) and CSL-Daily (Zhou_2021_CVPR) benchmarks as a gloss-based model, with BLEU scores of $29.03$ and $25.52$, respectively. The approach models keypoints extracted from the face, body, and both hands using an attention mechanism, and is jointly optimized with a CTC loss and a self-distillation loss. However, glosses require manual human annotation, which is labor-intensive and difficult to scale. In contrast, recent approaches attempt to model SLT at the sentence level and a gloss-free fashion. Building on this objective, recent works have incorporated LLMs into SLT frameworks (gong2024signllm; wong2024sign2gpt; chen-etal-2024-factorized), employed contrastive learning objectives (cheng2023cico), replaced gloss supervision with sentence-level embeddings (hamidullah2024slt), and extended SLT to multilingual scenarios (yin2022multislt; yazdani-etal-2025-continual; tan-etal-2025-multilingual; hamidullah-etal-2025-sonar).
@@ 2.3.  Hallucination
A model is said to hallucinate when it generates content that either lacks logical coherence or diverges from the source material, asserting information or events that are unfounded or inconsistent with the known facts (survey-of-hallucination). While hallucination detection has received increasing attention in the MT community, with studies on multilingual comparisons (guerreiro-etal-2023-hallucinations) and LLM-based approaches for both low- and high-resource languages (benkirane-etal-2024-machine), research on hallucination in SLT remains limited. zhang2023sltunet reported, based on manual analysis, that SLT models such as SLTUNET suffer greatly from hallucinations, where the generated translations often show limited correlation with the sign video. More recently, hamidullah2025grounding proposed a hallucination detection framework and further demonstrated that gloss-free SLT models are particularly prone to severe hallucinations compared to their gloss-based counterparts. These findings highlight the need for a deeper understanding of how evaluation metrics respond to hallucinated outputs and how robust current SLT models are under such conditions.
@@ 3.1.  Task and Datasets
In our experiments, we evaluate the limitations of current SLT metrics with respect to paraphrasing (including word reordering), hallucination sensitivity, and variation in translation length. We use the Phoenix-2014T  [^1] Phoenix-2014T was selected because it is the most widely used benchmark in SLT research and contains gloss-level annotations necessary for evaluating gloss-based models. dataset (Camgoz_2018_CVPR), covering weather forecasts in German sign language (DGS) to evaluate and compare the performance of various models in the above aspects.
@@
SLT models are prone to hallucinations, where the generated text diverges from the visual input. Yet, there has been no systematic investigation into how such hallucinations affect translation quality or the reliability of automatic SLT evaluation metrics. Motivated by findings from benkirane-etal-2024-machine, which demonstrate the effectiveness of LLMs for hallucination detection in MT, we employ Llama-3-70B to identify hallucinated outputs in a reference-based setting. [^2] The prompt used for hallucination detection is available in Appendix A.1. To analyze how evaluation metrics behave under varying degrees of hallucination, we consider a Severity Ranking scheme that categorizes outputs into four levels: No Hallucination, Small Hallucination, Partial Hallucination, and Full Hallucination.
@@ 3.2.  Evaluation Metrics
For our analysis, we consider the metrics as per muller-etal-2022-findings; muller-etal-2023-findings, specifically BLEU [^3] BLEU|nrefs:1|bs:25|tok:none|eff:no| case:mixed|smooth:exp|version:2.5.1 (via SacreBLEU; post-2018-call) for lexical overlap, ROUGE (lin-2004-rouge) [^4] ROUGE|metrics:rougeL|nrefs:1| stemmer:true|bs:25|version:0.1.2 for recall-oriented n-gram overlap, chrF (popovic-2015-chrf)555chrF|nrefs:1|bs:25|case:mixed|eff:yes| nc:6|nw:2|space:no|version:2.5.1 that uses character $n$-grams, and embedding-based BLEURT (sellam-etal-2020-bleurt) [^6] BLEURT v0.0.2 using checkpoint BLEURT-20. for semantic quality. We use bootstrap resampling for statistical significance reporting.
We also extend the traditional range of SLT evaluation metrics by presenting a sign language adapted version of G-Eval (liu-etal-2023-g). G-Eval introduces a framework that leverages prompting techniques to generate evaluation scores that strongly correlate with human preferences. The prompt is organized into three modules: (1) evaluation dimensions, which specify the aspects to be judged; (2) step-by-step reasoning, where Chain-of-Thought (CoT) (10.5555/3600270.3602070) guides the LLM through the evaluation procedure; and (3) scoring with references, which constrains the output format and incorporates human translations as ground truth. Following sato-etal-2024-tmu, we prompt the LLMs to evaluate the translation quality of the models in Section 3.3 for Adequacy and Fluency, each rated on a 5-point Likert scale (1155). Following fomicheva-etal-2022-mlqe; kocmi-federmann-2023-gemba, we further include GEMBA zero-shot Direct Assessment (kocmi-federmann-2023-gemba), where the LLM is prompted to evaluate each translation hypothesis and provide a quality score ranging from 0 (completely incorrect) to $100$ (perfect translation). Our LLM experimental setup includes four models: GPT-4.1-nano, Qwen3-8B, Llama-3.1-8B, and QWQ-32B. These were chosen to include both a proprietary LLM and also the open-source ones. [^7] All prompts used for G-Eval and GEMBA are provided in Appendices A.2 and A.3.
@@
SLT models are either gloss-based or gloss-free. We include three gloss-free models: SpaMo (hwang-etal-2025-spamo), SEM-SLT (hamidullah2024slt), and Signformer (yang2024signformerneededgeai). We also include TwoStream-SLT (10.5555/3600270.3601510) as a gloss-based model. We retrain these models to reproduce their reported results when the original model checkpoints are unavailable; however, our reproduced scores differ slightly from those originally reported.
@@
With the goal of edge AI, Signformer uses a combination of novel convolution, attention, and positional encoding to achieve competitive performance compared to gloss-based models while being quite smaller. Since the original paper did not specify how video features were extracted, we instead employ the S3D model pre-trained on both WLASL for ASL word recognition (li2020word) and the Kinetics-400 dataset for human action recognition (kay2017kineticshumanactionvideo), following the approach by (Chen2022ASM). We set learning rate to 0.001 and train it for a maximum of 100 epochs.
@@ 4.1.  Impact of Paraphrasing on SLT Evaluation
We present the results of LLM-based evaluators, G-Eval (Adequacy and Fluency) and GEMBA, comparing paraphrased and original translations on the Phoenix-2014T test set in Table 2. Most notably, paraphrasing generally improves G-Eval Adequacy and Fluency and GEMBA scores, consistent with the pattern observed for BLEURT. Llama3.1:8b is the only model that rates paraphrased translations as less fluent than the original ones. Among all evaluated LLMs, Llama3.1:8b tends to assign the highest Adequacy and Fluency scores, whereas gpt-4.1-nano assigns the lowest. We observe that LLM-based evaluators tend to prefer paraphrased translations over the original ones. This tendency likely arises because the paraphrases are themselves generated by an LLM, which may introduce stylistic patterns or linguistic preferences that align with the evaluator’s own training distribution, thereby biasing its judgments. Interestingly, this observation aligns with previous studies suggesting that LLMs exhibit a bias toward LLM-generated or paraphrased text (xu-etal-2024-pride; liu-etal-2023-g). To be able to make more general statements about evaluation metrics, we show the Pearson correlation between the lexical metrics, BLEURT, and GEMBA, computed using the average scores from Table 2 in Figure 1. We notice a high correlation among lexical overlap metrics, and GEMBA is also highly correlated with BLEURT, as both incorporate learned language model representations in their scoring. Our previous observations from Tables 1 and 2 are confirmed, with lexical metrics (BLEU, chrF, ROUGE) yielding consistent rankings, while LLM-based evaluations align closely with BLEURT, ranking TwoStream-SLT highest, followed by SpaMo, SEM-SLT, and Signformer.
@@ 4.2.  Impact of Hallucination on SLT Evaluation
We first analyze how hallucinations impact automatic metrics and then examine how robust SLT models are in the case of hallucination. Figure 2 illustrates the impact of hallucinations, categorized by their severity ranking, on various evaluation metrics across four SLT models. We omitted G-Eval in this setting because its results are reported on a Likert scale.  [^8] Typically ranging from 1 (strongly disagree/poor) to 5 (strongly agree/excellent) (likert1932technique). Moreover, we report hallucination detection results for GEMBA using gpt-4.1-nano. Notably, results show that there is a monotonic decline across all metrics, meaning that as hallucination severity level increases, translation quality drops. However, BLEU is an extremely sensitive metric; in cases such as Partial Hallucinations or Full Hallucinations, even minor rephrasings that remain semantically correct can cause a sharp decrease in BLEU scores. ROUGE and chrF decline more smoothly, since they account for partial overlaps at the subword or character level, respectively. GEMBA and BLEURT follow a similar trend, as they can still capture some semantic meaning even in extreme Full Hallucination cases where the surface wording differs substantially. However, we suspect that these metrics may under-penalize subtle hallucinations when the generated text is fluent, thereby overestimating the overall score. Similar observations regarding BLEURT’s overestimation and the data leakage issues in LLMs have been discussed by zeng-etal-2024-towards.
@@
Different from the main focus of the paper that concentrates on evaluating evaluation metrics, here we briefly take the opportunity to investigate the behavior of SLT models when hallucinations occurs. Overall, TwoStream-SLT and SpaMo demonstrate a consistent trend across all evaluation metrics, showing less hallucinations compared to other models. In contrast, Signformer exhibits notably lower average performance, aligning with the results presented in Table 1. To analyze the extent to which SLT models produce hallucinations, we present in Table 3 a breakdown of the number of hallucinated translations according to severity level. The results reveal that gloss-based TwoStream-SLT yields the lowest Full Hallucinations (24), indicating higher robustness to severe errors, whereas Signformer produces the highest number (177), confirming our earlier observations regarding its overall vulnerability to hallucination. This finding is in line with previous work showing that gloss-free models exhibit higher hallucination rates than gloss-based models (hamidullah2025grounding).